OpenStudy (ccchristian1):

The work of a student to solve a set of equations is shown: Equation A: y = 15 − 2z Equation B: 2y = 3 − 4z Step 1: −2(y) = −2(15 − 2z) [Equation A is multiplied by −2.] 2y = 3 − 4z [Equation B] Step 2: −2y = 15 − 2z [Equation A in Step 1 is simplified.] 2y = 3 − 4z [Equation B] Step 3: 0 = 18 − 6z [Equations in Step 2 are added.] Step 4: 6z = 18 Step 5: z = 3 In which step did the student first make an error? Step1 Step 2 Step 3 Step 4

OpenStudy (ccchristian1):

@TheSmartOne

OpenStudy (ccchristian1):

@AriPotta

OpenStudy (ccchristian1):

@campbell_st

OpenStudy (ccchristian1):

@563blackghost

TheSmartOne (thesmartone):

You're still taking the test?

OpenStudy (ccchristian1):

this is a different one

OpenStudy (ccchristian1):

08.09 Module Eight Test Part One

OpenStudy (ccchristian1):

@TheSmartOne

OpenStudy (campbell_st):

I'd say the teacher made the error in asking for a solution to a set of equations that have no solution... and then the next problem is asking to use elimination... the 2 lines are parallel and have no solution... y = 15 - 2z B: divide by 2 y = 3/2 - 2z I would have stopped there and said there is no solution.

OpenStudy (campbell_st):

I'd say the error by the student was attempting this inane question.

OpenStudy (ccchristian1):

so step 1 is wrong ?

OpenStudy (campbell_st):

the error is in the multiplication of equation A by -2 -2 * y = -2 * 15 - (-2)*2z it should be -2y = -30 + 4z

OpenStudy (ccchristian1):

okay